OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (anonymous):

since one of those equations is explicit in y, use substitution...

HanAkoSolo (jamierox4ev3r):

i would use substitution 5x+3(-2x+1)=8 5x-6x+3=8 -x+3=8 -3 -3 -x=5 x=-5

HanAkoSolo (jamierox4ev3r):

does that make sense @Viceroy ?

OpenStudy (anonymous):

It doesn't really make sense to me. Is that the answer?

HanAkoSolo (jamierox4ev3r):

that is the answer of the x variable, but you can also solve for y using x

HanAkoSolo (jamierox4ev3r):

y=-2(-5)+1 y=10+1 y=11 5(-5)+3y=8 -25+3y=8 +25 +25 3y=33 y=11 x=-5, y=11

HanAkoSolo (jamierox4ev3r):

then you could turn that into a slope intercept equation y=-5x+11

HanAkoSolo (jamierox4ev3r):

then, you could graph it if need. Does that make sense? if not, let me know what you want me to explain again

HanAkoSolo (jamierox4ev3r):

@Viceroy

OpenStudy (anonymous):

I get it. Thanks mate.